# Gaussian-Based Periodic Grand Canonical Density Functional Theory with Implicit Solvation for Computational Electrochemistry

### v2

@@ II.2.1 Solvated DFT energy
In general, the solvated DFT energy can be decomposed in the following form:Islam et al. (2023)
$E_{\text{DFT}}[\rho,\phi]=E_{\text{gas}}[\rho]+E_{\text{solv}}[\rho,\phi],$ (10)
@@
with $E_{\text{gas}}$ representing the energy in the absence of solvation, $E_{J,\text{solv}}$ being the Coulomb correction, $E_{\text{diel}}$ being the dielectric free energy, $E_{\text{ion}}$ being the ionic free energy, and $E_{\text{cd}}$ corresponding to nonelectrostatic cavitation and dispersion energies. There are two variational parameters: $\rho(\textbf{r})$ and $\phi(\textbf{r})$ which are the solute valence electron density and the overall electrostatic potential in the fluid environment, respectively. Here, $\rho$ represents the valence electronic density since the core is absent due to pseudopotentials. Later, we will use $\rho_{\text{sol}}$, $\rho_{\text{diel}}$, $\rho_{\text{ion}}$, $\rho_{\text{tot}}$, to represent the sum of nuclear and valence electronic density, bound solvent charge density, ionic charge density, and total charge density, respectively. $\phi_{\text{sol}}$ and $\phi\equiv\phi_{\text{tot}}$ are defined analogously for the electrostatic potential. We define $\phi_{\text{solv}}:=\phi_{\text{tot}}-\phi_{\text{sol}}$. The last four terms of Eq.˜11 comprise the solvation energy $E_{\text{solv}}$. The form of each energy component under linear fluid response is given in this section.
@@ II.2.2 Cavity functions
In continuum solvation frameworks, the solvent-accessible cavity function $s(\textbf{r})$ must be defined. This function is usually a twice-differentiable switching function that is zero in the solute and one in the continuum region. There are many possible formulations of the cavity functions.Letchworth-Weaver and Arias (2012); Andreussi, Dabo, and Marzari (2012); Fisicaro et al. (2017); Stein, Herbert, and Head-Gordon (2019) The cavity function defines the dielectric cavity $s_{\epsilon}(\textbf{r})$ which can vary between solvation models. The dielectric cavity defines the effective dielectric constant over the grid in these polarizable continuum models such that the dielectric constant is 1 in the solute region and is $\epsilon_{b}$, the bulk dielectric constant, in the solvent region.
@@
In our implementation, we followed the density-based cavity functions adopted by JDFTx and VASPsol.Gunceler et al. (2013); Mathew et al. (2019) These are known to be efficient and accurate cavities for most systems and are routinely used in computational electrocatalysis problems. In the local and linear polarizable continuum model (LPCM), the local density cavity function is parametrized by two tunable parameters, $\rho_{\text{cut}}$ and $\sigma$. We use $\rho_{\text{cut}}=0.00037$ $a_{0}^{-3}$ and $\sigma=0.6$.Gunceler et al. (2013)
@@
We have also implemented the charge-asymmetric nonlocally determined (CANDLE) shape function, which incorporates charge asymmetry into the shape function.Sundararaman and Goddard (2015) Nonlocality is incorporated into the CANDLE cavity through the convolution of the sphericalized density of a solvent molecule $w_{\text{lq}}$ and the valence electronic density $\rho(\textbf{r})$. This convolution defines an effective electronic density to build the cavity: $\overline{\rho}(\textbf{r})=(w_{\text{lq}}*\rho)\;(\textbf{r})$. Unique to CANDLE is the presence of a charge asymmetric term, which moves the solvation cavity closer to the solute for anions.Sundararaman and Goddard (2015) To obtain the dielectric cavity, the cavity function is smeared via convolution with a radial delta function so that the transition from solute to bulk solvent is less steep. The cavity function is parametrized by a solvent-dependent quantity $p_{\text{cav}}$, the coefficient on the charge asymmetric term. For water, $p_{\text{cav}}$ is 36.5. The quantity $Z_{\text{val}}$ is the effective solute nuclear charge, and $\hat{e}_{\nabla\overline{\rho}}$ is the unit vector parallel to $\nabla\overline{\rho}$.
@@ II.2.5 Ionic free energy
The ionic charge density is given by $\rho_{\text{ion}}\equiv-\frac{\epsilon_{b}\kappa^{2}}{4\pi}s_{\epsilon}(\textbf{r})$ where $\kappa$ is the inverse Debye screening length. The Debye screening length can be computed as a function of the bulk concentration of each ionic species $c_{i}$ via
@@ II.2.6 Nonelectrostatic energy
The final term in the free energy is associated with cavitation and dispersive forces. In most solvation models, the cavitation energy is approximated by scaling the surface area of the solvent accessible cavity with $\tau$, a parameter that resembles surface tension.Scherlis et al. (2006); Andreussi, Dabo, and Marzari (2012); Ringe et al. (2016) This is the form we use for our LPCM implementation
@@
In the CANDLE solvation model, a parameter-free nonlocally determined cavitation energy and solvent D2-type interactions comprise $E_{\text{cd}}$. Sundararaman, Gunceler, and Arias (2014); Grimme (2006)
@@ II.2.7 Implementation details
The equation for the solvation potential can be computed by taking the variation of free energy with respect to $\phi$. Upon differentiation, we obtain the linearized Poisson-Boltzmann equation, which can be solved efficiently in reciprocal space using the preconditioned conjugate gradient algorithm.Letchworth-Weaver and Arias (2012); Mathew et al. (2019)
$\nabla(\epsilon(\textbf{r})\cdot\nabla\phi(\textbf{r}))+\epsilon_{b}\kappa^{2}s_{\epsilon}(\textbf{r})\phi(\textbf{r})=4\pi\rho_{\text{sol}}(\textbf{r}).$ (23)
@@
$V_{\text{solv}}=\phi_{\text{solv}}-\left(\frac{\epsilon_{\text{b}}-1}{8\pi}\right)\frac{\delta s_{\epsilon}}{\delta\rho}|\nabla\phi|^{2}-\beta c_{\text{ion}}z_{\text{ion}}^{2}\frac{\delta s_{\epsilon}}{\delta\rho}\phi+\frac{\delta E_{\text{cd}}}{\delta\rho}.$ (24)
At each SCF iteration, we solve for the updated total electrostatic potential $\phi$ based on the new density. This procedure does not change whether we are performing a constant electron number or a constant potential calculation. The general workflow of our method is detailed in Algorithm 1. For completeness, we provide the full derivation of $\delta E/\delta\rho$ in Appendix A.
@@
  Solve Eq.˜23 for $\phi$.
  Solve for $V_{\text{solv}}$ using Eq.˜24 using $\phi$.
@@ III Computational Details
Both the grand canonical DFT method and implicit solvation models are implemented in a development version of Q–Chem with Periodic Boundary Conditions (QCPBC).Epifanovsky et al. (2021); Lee et al. (2021, 2022); Rettig, Lee, and Head-Gordon (2023) The code uses the Goedecker-Teter-Hutter (GTH) pseudopotentials optimized for the Perdew-Burke-Ernzerhof (PBE) functional and the corresponding optimized basis sets for the GTH pseudopotentials.Perdew, Burke, and Ernzerhof (1996); Goedecker, Teter, and Hutter (1996); VandeVondele and Hutter (2007) Unless otherwise noted, the TZVP-MOLOPT-PBE-GTH basis set developed by CP2K was used for all calculations.Kühne et al. (2020) The Gaussian and Plane Waves method (GPW) was used for the evaluation of two-body integrals.VandeVondele et al. (2005) Unless otherwise specified, all metallic slab calculations were performed with a five-layer slab with one atom per layer with 15 Å of vacuum. A $12\times 12\times 1$ $\Gamma$-centered $\mathbf{k}$-point mesh was used for such calculations. Fermi smearing with a smearing temperature of 0.001 a.u. was employed for all metallic calculations.
@@ IV.1 Verification of Implemented Solvation Models
Table 1: Comparison of the accuracy of LPCM and CANDLE between JDFTx (plane-waves basis set with ultrasoft pseudopotentials)Sundararaman et al. (2017) and QCPBC (our implementation). Experimental values for JDFTx LPCMSundararaman and Schwarz (2017) and JDFTx CANDLESundararaman and Goddard (2015) are obtained from corresponding references.
To confirm the transferability of solvation models developed for plane wave electronic structure codes and different pseudopotentials to our GTO-based code, we computed single-point solvation energies for a benchmark set of 240 neutral molecules, 52 cations, and 60 anions.National Institute of Standards and Technology (2022) This is also the dataset used for the calibration of previous solvation models.Andreussi, Dabo, and Marzari (2012); Gunceler et al. (2013); Dupont, Andreussi, and Marzari (2013) For all calculations, we used the PBE functional, and all molecules were placed in a 20 Å cubic box with the GPW plane wave cutoff as 3000 eV. The results are shown in Table 1. The results suggest that these solvation energies do not have a strong dependence on basis set choice or pseudopotential.
The presence of electrolytes in our solvation model sets a natural reference electrostatic potential in the bulk electrolyte. Still, in most periodic electronic structure codes, the electrostatic potential is chosen such that the integral of the electrostatic potential is 0. Our implementation automatically shifts the reference potential, so one does not need to add a shifting quantity to the Fermi energy, such as EFERMI_SHIFT in a VASPsol calculation.Mathew et al. (2019) We verify our reference potential by plotting the grand canonical free energy near the PZC of an Ag(100) slab. We observe the expected parabolic shape to the energy curve as well as a free energy maximum at neutral charge.Santos and Schmickler (2004) The details of this are included in Appendix B.
One potential weakness of implicit solvation models is the undercharging of electrodes as the electrode potential varies. Most such polarizable continuum models are fit to solvation energies of small molecules and ions. As a result, the capacitance of most metal electrodes is underestimated.Sundararaman and Schwarz (2017) Reparameterizations of preexisting solvation models have been presented to mitigate this issue, often at the cost of accuracy in energy evaluation.Sundararaman and Schwarz (2017) To combat this, we have implemented the CANDLE solvation model, which largely resolves this issue for negatively charged slabs by constructing a smaller solute cavity for the anionic regions of the solute. With our implementation of CANDLE, we were able to reproduce the results presented by Sundararaman and coworkers, where CANDLE accurately describes the charging behavior of negatively charged silver slabs.Sundararaman and Schwarz (2017) The results are presented in Fig. 1. It should be noted that the capacitance is still underestimated for positively charged electrodes even with CANDLE.
@@
Figure 2: Energy ($\Delta E$) and charge ($\Delta Q$) convergence plot with CANDLE for two systems: (a) Cu(111) at 1 V SHE and (b) Pt(111) at 0 V SHE. The convergence is compared with results from JDFTx, which implements the auxiliary Hamiltonian method for minimization of the grand canonical free energy.Sundararaman, Goddard, and Arias (2017)
@@ IV.2 Grand canonical SCF convergence
We compared the convergence behavior of our direct minimization approach to the auxiliary Hamiltonian method implemented in JDFTx, an efficient variational optimization approach for plane wave-based codes.Sundararaman, Goddard, and Arias (2017) In particular, we compared the SCF convergence against JDFTx for the two benchmark systems described in Sundararaman et. al.: Cu(111) at 1 V and Pt(111) at 0 V relative to the computational standard hydrogen electrode (SHE).Sundararaman, Goddard, and Arias (2017); Nørskov et al. (2004) Consistent with their work, we use the CANDLE solvation model for this set of calculations. The results are shown in Fig.˜2. Our method demonstrates significantly better convergence over JDFTx, as we can converge energies and charge in about half the number of conjugate gradient iterations in both cases. This provides evidence that our direct density matrix minimization approach is a robust strategy for grand canonical SCF calculations.
@@ IV.3 Computational efficiency
We also determine the computational overhead for including solvation in calculations. Ideally, one would like this overhead to be negligible compared to the rest of the computational components that are also common to canonical DFT. Because within each SCF iteration, we solve Eq.˜23 via the conjugate-gradient algorithm, we want to confirm that this iterative procedure does not significantly increase the runtime relative to the corresponding gas-phase calculation.
@@ IV.4 Computational standard hydrogen electrode reference potential
To perform constant potential calculations, an absolute electrode potential relative to vacuum must be established. In this work, we computed the absolute potential of the SHE relative to the vacuum using the method described by JDFTx.Letchworth-Weaver and Arias (2012); Gunceler et al. (2013) Experimental potentials of different facets of copper, silver, and gold are well benchmarked, so the computational potentials of zero charge (PZC) are fitted to these values to obtain the computational hydrogen electrode reference.Trasatti and Lust (1999) These metals are chosen because they interact minimally with the electrolyte. For all calculations, 1.0 M of non-interacting electrolyte was assumed. The fitting obtained using our method is shown in Fig. 4. The computed SHE values using the PBE functional for the LPCM and CANDLE models are 4.65-4.65 V and 4.58-4.58 V, respectively. This value is in close agreement with values reported from other implementations of this method in VASP (LPCM: 4.6-4.6 V) and JDFTx (LPCM: 4.68-4.68 V, CANDLE: 4.66-4.66 V).Gunceler et al. (2013); Mathew et al. (2019) We also fit SHE values for the RPBE functional that has been shown to give improved energetic results for absorption energies and other surface-related reactions.Hammer, Hansen, and Nørskov (1999) We obtained computed SHE values of 4.48-4.48 V and 4.44-4.44 V for LPCM and CANDLE models respectively.

### v3

@@ II.2.1 Solvated DFT energy
In general, the solvated DFT energy can be decomposed in the following form:52
$E_{\text{DFT}}[\rho,\phi]=E_{T}+E_{J}[\rho]+E_{\text{xc},\text{gas}}[\rho]+E_{\text{solv}}[\rho,\phi],$ (10)
@@
with $E_{\text{gas}}$ representing the energy in the absence of solvation, $E_{J,\text{solv}}$ being the Coulomb correction, $E_{\text{diel}}$ being the dielectric free energy, $E_{\text{ion}}$ being the ionic free energy, and $E_{\text{cd}}$ corresponding to nonelectrostatic cavitation and dispersion energies. There are two variational parameters: $\rho(\textbf{r})$ and $\phi(\textbf{r})$ which are the solute valence electron density and the overall electrostatic potential in the fluid environment, respectively. Here, $\rho$ represents the valence electronic density since the core is absent due to pseudopotentials. Later, we will use $\rho_{\text{sol}}$, $\rho_{\text{diel}}$, $\rho_{\text{ion}}$, $\rho_{\text{tot}}$, to represent the sum of nuclear and valence electronic density, bound solvent charge density, ionic charge density, and total charge density, respectively. $\phi_{\text{sol}}$ and $\phi\equiv\phi_{\text{tot}}$ are defined analogously for the electrostatic potential. We define $\phi_{\text{solv}}:=\phi_{\text{tot}}-\phi_{\text{sol}}$. The form of each energy component of $E_{\text{solv}}$ under linear fluid response is given in this section.
@@ II.2.2 Cavity functions
In continuum solvation frameworks, the solvent-accessible cavity function $s(\textbf{r})$ must be defined. This function is usually a twice-differentiable switching function that is zero in the solute and one in the continuum region. There are many possible formulations of the cavity functions.11, 44, 53, 54 The cavity function defines the dielectric cavity $s_{\epsilon}(\textbf{r})$ which can vary between solvation models. The dielectric cavity defines the effective dielectric constant over the grid in these polarizable continuum models such that the dielectric constant is 1 in the solute region and is $\epsilon_{b}$, the bulk dielectric constant, in the solvent region.
@@
In our implementation, we followed the density-based cavity functions adopted by JDFTx and VASPsol.42, 9 These are known to be efficient and accurate cavities for most systems and are routinely used in computational electrocatalysis problems. In the local and linear polarizable continuum model (LPCM), the local density cavity function is parametrized by two tunable parameters, $\rho_{\text{cut}}$ and $\sigma$. We use $\rho_{\text{cut}}=0.00037$ $a_{0}^{-3}$ and $\sigma=0.6$.42
@@
We have also implemented the charge-asymmetric nonlocally determined (CANDLE) shape function, which incorporates charge asymmetry into the shape function.43 Nonlocality is incorporated into the CANDLE cavity through the convolution of the sphericalized density of a solvent molecule $w_{\text{lq}}$ and the valence electronic density $\rho(\textbf{r})$. This convolution defines an effective electronic density to build the cavity: $\overline{\rho}(\textbf{r})=(w_{\text{lq}}*\rho)\;(\textbf{r})$. Unique to CANDLE is the presence of a charge asymmetric term, which moves the solvation cavity closer to the solute for anions.43 To obtain the dielectric cavity, the cavity function is smeared via convolution with a radial delta function so that the transition from solute to bulk solvent is less steep. The cavity function is parametrized by a solvent-dependent quantity $p_{\text{cav}}$, the coefficient on the charge asymmetric term. For water, $p_{\text{cav}}$ is 36.5. The quantity $Z_{\text{val}}$ is the effective solute nuclear charge, and $\hat{e}_{\nabla\overline{\rho}}$ is the unit vector parallel to $\nabla\overline{\rho}$.
@@ II.2.5 Ionic free energy
The ionic charge density is given by $\rho_{\text{ion}}\equiv-\frac{\epsilon_{b}\kappa^{2}}{4\pi}s_{\epsilon}(\textbf{r})\phi(\mathbf{r})$ where $\kappa$ is the inverse Debye screening length. The Debye screening length can be computed as a function of the bulk concentration of each ionic species $c_{i}$ via
@@ II.2.6 Nonelectrostatic energy
The final term in the free energy is associated with cavitation and dispersive forces. In most solvation models, the cavitation energy is approximated by scaling the surface area of the solvent accessible cavity with $\tau$, a parameter that resembles surface tension.55, 44, 56 This is the form we use for our LPCM implementation, where we have $\tau=5.4\times 10^{-6}\;E_{h}\cdot a_{0}^{-2}$:
@@
In the CANDLE solvation model, a parameter-free nonlocally determined cavitation energy and solvent D2-type interactions comprise $E_{\text{cd}}$. 57, 58
@@ II.2.7 Implementation details
The equation for the solvation potential can be computed by taking the variation of free energy with respect to $\phi$. Upon differentiation, we obtain the linearized Poisson-Boltzmann equation, which can be solved efficiently in reciprocal space using the preconditioned conjugate gradient algorithm.11, 9
$\nabla(\epsilon(\textbf{r})\cdot\nabla\phi(\textbf{r}))+\epsilon_{\text{b}}\kappa^{2}s_{\epsilon}(\textbf{r})\phi(\textbf{r})=4\pi\rho_{\text{sol}}(\textbf{r}).$ (23)
@@
$V_{\text{solv}}=\phi_{\text{solv}}-\left(\frac{\epsilon_{\text{b}}-1}{8\pi}\right)\frac{\delta s_{\epsilon}}{\delta\rho}|\nabla\phi|^{2}-\beta c_{\text{ion}}z_{\text{ion}}^{2}\frac{\delta s_{\epsilon}}{\delta\rho}\phi^{2}+\frac{\delta E_{\text{cd}}}{\delta\rho}.$ (24)
At each SCF iteration, we solve for the updated total electrostatic potential $\phi$ based on the new density. This procedure does not change whether we are performing a constant electron number or a constant potential calculation. Our implementation of the implicit solvent method requires little additional memory, since only the electronic density, its derived quantities, and the electrostatic potential need to be stored on a uniform grid. The density and its gradients are already evaluated on this grid for the computation of exchange-correlation contributions. The general workflow of our method is detailed in Algorithm 1. For completeness, we provide the full derivation of $\delta E/\delta\rho$ in the Supporting Information.
@@
  Solve Eq. 23 for $\phi$.
  Solve for $V_{\text{solv}}$ using Eq. 24 using $\phi$.
@@ III Computational Details
Both the grand canonical DFT method and implicit solvation models are implemented in a development version of Q–Chem with Periodic Boundary Conditions (QCPBC).59, 60, 61, 62 Our implementation supports shared-memory parallel execution across multiple CPU cores. The code uses the Goedecker-Teter-Hutter (GTH) pseudopotentials optimized for the Perdew-Burke-Ernzerhof (PBE) functional and the corresponding optimized basis sets for the GTH pseudopotentials.63, 64, 65 Unless otherwise noted, the TZVP-MOLOPT-PBE-GTH basis set developed by CP2K was used for all calculations.45 The Gaussian and Plane Waves method (GPW) was used for the evaluation of two-body integrals.66 Unless otherwise specified, all metallic slab calculations were performed with a five-layer slab with one atom per layer with 15 Å of vacuum. A $12\times 12\times 1$ $\Gamma$-centered $\mathbf{k}$-point mesh was used for such calculations. Fermi smearing with a smearing temperature of 0.001 a.u. was employed for all metallic calculations. All calculations were performed on an AMD EPYC 9654 Processor where 16 cores and 64 GB of memory were allocated.
@@ IV.1 Verification of Implemented Solvation Models
Table 1: Comparison of the accuracy of LPCM and CANDLE between JDFTx (plane waves basis set with ultrasoft pseudopotentials)31 and QCPBC (our implementation). Errors for JDFTx LPCM67 and JDFTx CANDLE43 are obtained from corresponding references.
To confirm the transferability of solvation models developed for plane wave electronic structure codes and different pseudopotentials to our GTO-based code, we computed single-point solvation energies for a benchmark set of 240 neutral molecules, 52 cations, and 60 anions.68 This is also the dataset used for the calibration of previous solvation models.44, 42, 69 For all calculations, we used the PBE functional, and all molecules were placed in a 20 Å cubic box with the GPW plane wave cutoff as 3000 eV. The results are shown in Table 1. The results suggest that these solvation energies do not have a strong dependence on basis set choice or pseudopotential.
The presence of electrolytes in our solvation model sets a natural reference electrostatic potential in the bulk electrolyte. Still, in most periodic electronic structure codes, the electrostatic potential is chosen such that the integral of the electrostatic potential is 0. Our implementation automatically shifts the reference potential, so one does not need to add a shifting quantity to the Fermi energy, such as EFERMI_SHIFT in a VASPsol calculation.9 We verify our reference potential by plotting the grand canonical free energy near the PZC of an Ag(100) slab. We observe the expected parabolic shape to the energy curve as well as a free energy maximum at neutral charge.70 The details of this are included in the Supporting Information.
One potential weakness of implicit solvation models is the undercharging of electrodes as the electrode potential varies. Most such polarizable continuum models are fit to solvation energies of small molecules and ions. As a result, the capacitance of most metal electrodes is underestimated.67 Reparameterizations of preexisting solvation models have been presented to mitigate this issue, often at the cost of accuracy in energy evaluation.67 To combat this, we have implemented the CANDLE solvation model, which largely resolves this issue for negatively charged slabs by constructing a smaller solute cavity for the anionic regions of the solute. With our implementation of CANDLE, we were able to reproduce the results presented by Sundararaman and coworkers, where CANDLE accurately describes the charging behavior of negatively charged silver slabs.67 The results are presented in Fig. 1. It should be noted that the capacitance is still underestimated for positively charged electrodes even with CANDLE.
@@
Figure 2: Energy ($\Delta E$) and charge ($\Delta Q$) convergence plot with CANDLE for two systems: (a) Cu(111) at 1 V SHE and (b) Pt(111) at 0 V SHE. The convergence is compared with results from JDFTx, which implements the auxiliary Hamiltonian method for minimization of the grand canonical free energy.17
@@ IV.2 Grand canonical SCF convergence
We compared the convergence behavior of our direct minimization approach to the auxiliary Hamiltonian method implemented in JDFTx, an efficient variational optimization approach for plane wave-based codes.17 In particular, we compared the SCF convergence against JDFTx for the two benchmark systems described in Sundararaman et. al.: Cu(111) at 1 V and Pt(111) at 0 V relative to the computed standard hydrogen electrode (SHE).17, 71 Consistent with their work, we use the CANDLE solvation model for this set of calculations. The results are shown in Fig. 2. Our method demonstrates significantly better convergence over JDFTx, as we can converge energies and charge in about half the number of conjugate gradient iterations in both cases. This provides evidence that our direct density matrix minimization approach is a robust strategy for grand canonical SCF calculations.
@@ IV.3 Computational efficiency
We also determine the computational overhead for including solvation in calculations. Ideally, one would like this overhead to be negligible compared to the rest of the computational components that are also common to canonical DFT. Because within each SCF iteration, we solve Eq. 23 via the conjugate-gradient algorithm, we want to confirm that this iterative procedure does not significantly increase the runtime relative to the corresponding gas-phase calculation.
@@
A direct comparison of overall timings between GTO- and plane wave-based codes can be challenging due to the differences in pseudopotentials, grid density, and basis set errors. Our solvation solver and related computational kernels are identical to those previously developed for plane waves.11, 9 Hence, the associated overhead for solvation is the same in our code and plane wave codes. The primary computational bottleneck in these calculations is the formation of the Fock matrix, which lies beyond the scope of this work, but we have recently reported progress.72 As a test, we also ran five-layer Cu(100) and Ag(100) slabs using our LPCM implementation and JDFTx on 32 threads on an AMD Ryzen 9 7950X CPU with the PBE functional. For the JDFTx calculations, we employed the recommended GBRV pseudopotentials with the associated energy cutoff of 20 $E_{h}$. In both cases, the calculations were initialized from a converged gas-phase electronic density. Our implementation required 79 s and 103 s to converge Cu(100) and Ag(100), respectively, while JDFTx required 77 s and 75 s. Although these timings are not directly comparable, they indicate that our GTO-based implementation achieves performance similar to existing plane wave implementations.
@@ IV.4 Computational standard hydrogen electrode reference potential
To perform constant potential calculations, an absolute electrode potential relative to vacuum must be established. In this work, we computed the absolute potential of the SHE relative to the vacuum using the method described by JDFTx.11, 42 Experimental potentials of different facets of copper, silver, and gold are well benchmarked, so the computational potentials of zero charge (PZC) are fitted to these values to obtain the computational hydrogen electrode reference.73 These metals are chosen because they interact minimally with the electrolyte. For all calculations, 1.0 M of non-interacting electrolyte was assumed. The fitting obtained using our method is shown in Fig. 4. The computed SHE values using the PBE functional for the LPCM and CANDLE models are 4.65-4.65 V and 4.58-4.58 V, respectively. This value is in close agreement with values reported from other implementations of this method in VASP (LPCM: 4.6-4.6 V) and JDFTx (LPCM: 4.68-4.68 V, CANDLE: 4.66-4.66 V).42, 9 We also fit SHE values for the RPBE functional that has been shown to give improved energetic results for absorption energies and other surface-related reactions.74 We obtained computed SHE values of 4.48-4.48 V and 4.44-4.44 V for LPCM and CANDLE models respectively.